OpenStudy (anonymous):

i need to know how to solve this integral without a calculator. i assume its by doing it graphically but i am not sure how to do that. 2x^3+3 going from 0 to 2

OpenStudy (anonymous):

With that you mean, you're not allowed to compute the integral at all and plugin the values? \[\large \int_0^2 (2x^3+3)dx= \left.\frac{1}{2}x^4+3x \right|_{x=0}^{x=2}=14\] No calculator needed for that, although I confess \(2^4\) can be a bit of a struggle in the head, at least for me hehe.

OpenStudy (loser66):

hahaha... me too!!

OpenStudy (anonymous):

ah ok i get it now thanks